Let the joint pdf of X , Y be , f (x, y) = 2 , 0 ≤ x ≤ y ≤ 1 and




f1(x) = 2(1 − x) , 0 ≤ x ≤ 1. Find E[X|Y]

1
Expert's answer
2021-12-21T17:04:01-0500

"E[X|Y]=\int xf(x|y)dx"


"f(x|y)=\frac{f(x,y)}{f(y)}=f(x)"


"E[X|Y]=\int xf(x)dx"


"E[X|Y]=2\int^1_0x(1-x)dx=(x^2-2x^3/3)|^1_0=1-2/3=1/3"